OpenStudy (anonymous):

Newton currently has an account balance of $6,225.18. He opened the account 15 years ago with a deposit of $4,543.11. If the interest compounds daily, what is the interest rate on the account? 30.3% 7.7% 58.3% 2.1%

OpenStudy (anonymous):

thank you satellite

OpenStudy (anonymous):

\[6,225.18=4543.11\times (1+\frac{r}{365})^{365\times 15}\]

OpenStudy (anonymous):

Your the best thank you=D

OpenStudy (anonymous):

yw do you know how to solve?

OpenStudy (anonymous):

Yep=D

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I got 2.1% is that correct?